OpenStudy (anonymous):

What is the factored form of the expression? k^2 – 9h^2 A. (k + 3h)(k + 3h) B. h^2(k + 3)(k – 3) C. (k + 3h)(k – 3h) D. (k – 3h^2)(k + 3) A, B, C, or D? Please explain (:

OpenStudy (anonymous):

first we need to correct the form of 9 which should be 3^2 now since everything is in squared form we can apply the formula a^2+b^2 = (a-b)(a+b) k^2- (3h)^2 (k-3h)(k+3h) C

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Thanks a lot guys (:

OpenStudy (anonymous):

k^2 - 9h^2 k^2 - (3h)^2 =(k - 3h) (k+3h) from the formula a^2-b^2 = (a-b)(a+b)